y + 4 over y=-5
show work

Answer by stanbon(75887)

y + 4 over y=-5
y+(4/y)=5
Multiply both sides by y to get:
y^2+4=5y
y^2-5y+4=0
(y-4)(y-1)=0
y=4 or y=1
Cheers,
stan H.